a stone is dropped into a well it travels in air for 5 seconds and touches the water available in well approximate and find the depth of well


1
Expert's answer
2021-08-18T11:07:34-0400

The depth of well

"d=\frac{gt^2}{2}=\frac{9.8\:\rm m/s^2*(5\: s)^2}{2}=122.5\:\rm m"